OpenStudy (darkbluechocobo):

f(x) = x^2 + 2x + 3 Help finding the vertex

OpenStudy (xapproachesinfinity):

do some massage to that quadratic function complete the square so you can write it in standard form \(\rm f(x)=a(x-h)^2+k\)

OpenStudy (xapproachesinfinity):

or just use \(h=-b/2a\) so \(k=f(-b/2a)\) (h,k) is the vertex

OpenStudy (darkbluechocobo):

What would be h and k in this would it be 2 and 3?

OpenStudy (darkbluechocobo):

Alright

OpenStudy (darkbluechocobo):

@xapproachesinfinity is this (-1,-4)